OpenStudy (anonymous):

The sum of two numbers is 56, and their difference is 10. What are the numbers? A) 33 and 23 B) 30 and 26 C) 20 and 36 D) 20 and 30 I am guessing A, but not sure. @kelliegirl33

OpenStudy (therealmeeeee):

let me see

OpenStudy (anonymous):

Let one number by x and the other number be y. Sum means add. Difference means subtract. So they are telling you that x + y = 56. And your second equation is x - y = 10. You have two equations and two unknowns. Do you how to solve this? I can explain if not.

OpenStudy (therealmeeeee):

yes your right i got the same answer:)

OpenStudy (anonymous):

Okay, thanks TheRealMeeeee (:.

OpenStudy (anonymous):

You don't have to guess. It's easy to solve. In the first equation, isolate x so that x = 56 - y. Then substitute that in for x in the second equation.

OpenStudy (therealmeeeee):

np

OpenStudy (anonymous):

I know, just wanted to make sure my answer was correct. Thanks anyways jam333.